# Inflationary Dynamics in Einstein-Gauss-Bonnet Gravity Using New Slow-Roll Approximations considering Generalised Reheating

### v1

@@ V.2 New approximation II
The second way to get $\delta_{1}(\phi)$ is the following. We neglect the term proportional to $\delta_{1}^{2}$ in Eq. (5.33) and get a nonzero solution:
@@
Using Eq. (4.19), we get
$\frac{dQ}{dN}=\frac{V_{,\phi}\,\delta_{1}}{12\xi_{,\phi}Q(1-\delta_{1})}+\frac{V\delta_{1}\delta_{2}}{6U_{0}(1-\delta_{1})^{2}}=\frac{U_{0}\,V_{,\phi}\,\delta_{1}}{2\xi_{,\phi}V}+\frac{V\delta_{1}\delta_{2}}{6U_{0}(1-\delta_{1})^{2}}\,.$ (5.47)
So,
$\varepsilon_{1}={}-\frac{3\,U^{2}_{0}\,V_{,\phi}}{2V^{2}\xi_{,\phi}}\delta_{1}(1-\delta_{1})-\frac{\delta_{1}\delta_{2}}{2\,(1-\delta_{1})}\,.$ (5.48)
From 4.20, we get
$\dot{\psi}\approx\frac{U_{0}\delta_{1}}{2\xi_{\phi}}\left(\delta_{2}+\varepsilon_{1}-\frac{3U_{0}^{2}\xi_{,\phi\phi}\delta_{1}}{V\,\xi_{,\phi}^{2}}\right)\,.$ (5.49)
Substituting formulae 5.46, 5.48, and 5.49 into 2.5, we get
$\frac{U_{0}\delta_{1}}{2\xi_{\phi}}\left(\delta_{2}+\varepsilon_{1}-\frac{3U_{0}^{2}\xi_{,\phi\phi}\delta_{1}}{V\,\xi_{,\phi}^{2}}\right)+\frac{3U_{0}\delta_{1}}{2\xi_{,\phi}}+V_{,\phi}+\frac{12\xi_{,\phi}V^{2}\,V_{,\phi}}{36\,U_{0}^{2}(1-\delta_{1})^{2}}(1-\varepsilon_{1})=0$ (5.50)
Multiplying this equation to $(1-\delta_{1})^{2}$ and supposing that any products of slow-roll parameters are negligible, we obtain the following linear equation in $\delta_{1}$:
$\frac{9}{2}\left(\frac{U_{0}^{3}}{\xi_{,\phi}}-V_{,\phi}\,U_{0}^{2}\right){\delta_{1}}+3\,V_{,\phi}\,U_{0}^{2}+V^{2}\xi_{,\phi}=0.$ (5.51)
Solving this equation, we get
@@
Now we can express $Q$, $\chi$, $N_{,\phi}\,$, and $\epsilon_{1}$ via $\phi$:
$Q(\phi)\simeq\frac{V}{6U_{0}(1-\delta_{1}(\phi))}=\frac{3{U_{0}}V(U_{0}-\xi_{,\phi}V_{,\phi})}{2(9U_{0}^{3}-3U_{0}^{2}\xi_{,\phi}V_{,\phi}+2\xi_{,\phi}^{2}V^{2})},$ (5.53)
$\chi=\frac{U_{0}\delta_{1}}{2\xi_{,\phi}Q}\simeq{}-\frac{2(3U_{0}^{2}V_{,\phi}+\xi_{,\phi}V^{2})(9U_{0}^{3}-3U_{0}^{2}\xi_{,\phi}V_{,\phi}+2\xi_{,\phi}^{2}V^{2})}{27U_{0}^{2}V{(U_{0}-\xi_{,\phi}V_{,\phi})}^{2}},$ (5.54)
@@ VI Inflationary analysis
In the subsequent sections, we will provide the calculation of inflationary observable for PLP in the new slow roll approximations. For our analysis we choose the following form of the EGB coupling [78, 79]
@@
where, $\xi_{1}$, $\xi_{2}$ are the coupling constant and $V_{0}$ is the scale of inflation which we fix from scalar power spectrum. Substituting 3.17 and 6.58 in 5.32, we can derive the explicit form of the effective potential (keeping $M_{p}=1$):
@@
Figure 6.1: Plots of $r-n_{s}$ obtained by using Slow-Roll Approximation I for different values of $n$ and $q$ we vary N from $50$ to $60$. The red color line denotes the $n=1$, the dashed green line is for $n=2$, the dotted blue line is for $n=3$ and the black dot-dashed line signifies the $n=4$. Where the small and large dots of different color signify the $\text{N}=50$ and $\text{N}=60$ respectively. The light blue shaded area is for the $2-\sigma$ bounds, while the light green shaded area represents the $1-\sigma$ bounds on $n_{s}$ from Planck’18 [13]
VI.1 Inflationary Observables in Slow-Roll I
Figure 6.2: Plots of $r-n_{s}$ obtained by using Slow-Roll Approximation II for different values of $n$ and $q$ we vary N from $50$ top $60$. The red color line denotes the $n=1$, the dashed green line is for $n=2$, the dotted blue line is for $n=3$ and the black dot-dashed line signifies the $n=4$. Where the small and large dots of different colors signifies the $\text{N}=50$ and $\text{N}=60$ respectively. The light blue shaded area is for the $2-\sigma$ bounds, while the light green shaded area represents the $1-\sigma$ bounds on $n_{s}$ from Planck’18 [13]
Of course, the results look cumbersome, to say the least, but one should appreciate the fact that using these techniques, one can still get analytical solutions.
Now, we will discuss the PLP inflation in slow roll approximation I. To calculate the inflationary observables first we calculate the slow roll parameters, using 5.44 and 5.40 we compute $\varepsilon_{1},\delta_{1}$. It is then straightforward to obtain $\varepsilon_{2}$ and $\delta_{2}$ using 5.45. When studying inflation another important quantity is the number of e-folds $N$. Thus, from 5.43, one can write $N$ for a given $\phi$. To solve the 5.43 we followed the same prescription as mentioned in [1], and we have taken $N=60$ at the end of inflation whereas $N=0$ corresponds to the onset of the inflation.
We set the initial field value for solving 5.43 at the end of inflation which can be calculated from $\varepsilon_{1}=1$. It is difficult to solve the 5.43 analytically, we then employ the numerical techniques. Equipped with all the preliminary equations in hand, one can finally compute the scalar spectral index ($n_{s}$), the tensor to the scalar ratio ($r$), and the amplitude of scalar power spectrum ($A_{s}$) using 4.27,4.28 and 4.29 respectively. In our analysis we take $\xi_{1}=10$ and $\xi_{2}=0.4$, these values of the coupling parameters are chosen so that the constraints on $n_{s}$ and $r$ does not contradict with the current CMB bounds. Whereas, $V_{0}$ can be fixed by matching $A_{s}=2.09\times 10^{-9}$ at the pivot scale.
As we have previously mentioned $n$ and $q$ vary from $1$ to $4$, which results in the $16$ different class of potential. We have analyzed for all the possible combinations. In fig. 6.1 we present the results on the $r-n_{s}$ where Slow-Roll I approximation has been utilized. It is evident from fig. 6.1 that almost all the obtained values of $r$ and $n_{s}$ for combination of potential parameters $n$ and $q$ are well inside the constraints imposed by the Planck 18. Explicit numerical values of the inflationary observables can be found in table 6.1.
@@ VI.1 Inflationary Observables in Slow-Roll I
$N$ $r$ $n_{s}$ $r$ $n_{s}$ $r$ $n_{s}$ $r$ $n_{s}$
$50$ 0.0188761    0.969662 0.00750292   0.961960 0.0058097   0.958572 0.0056763   0.956782
$60$ 0.0146428   0.975524 0.00533771   0.968838 0.003992   0.965786 0.00386787   0.965061
VI.2 Inflationary Observables in Slow-Roll II
Using the obtained expressions of $\delta_{1}(\phi)$ and $\varepsilon_{1}(\phi)$ and Eq. 5.45, one can now calculate parameters $\varepsilon_{2}(\phi)$ and $\delta_{2}(\phi)$. One can again get the analytic expressions for the slow-roll parameters using the new approximation-II as given in A.2. After this, one can in the same spirit obtain $n_{s}(\phi)$, $r(\phi)$, and $A_{s}(\phi)$ using (4.27)–(4.29).
In this section we will discuss the PLP inflation in slow roll approximation II. We employ the same method to calculate the inflationary observables as we have employed in Slow Roll Approximation I. First we calculate the slow roll parameters ($\varepsilon_{1},\varepsilon_{2},\delta_{1},\delta_{2}$), using 5.56 and 5.52 we compute $\varepsilon_{1},\delta_{1}$. After this, it is straightforward to obtain $\varepsilon_{2}$ and $\delta_{2}$ using 5.45. Using 5.55 we calculate the number of e-folds $N$ by adopting the same technique as we did in the previous subsection(Slow-Roll Approximation I). Equipped with all the preliminaries we can finally compute the scalar spectral index ($n_{s}$), tensor to scalar ratio ($r$), and amplitude of scalar power spectrum ($A_{s}$) using 4.27,4.28 and 4.29 respectively. Here also, we take $\xi_{1}=10$ and $\xi_{2}=0.4$, these values of the coupling parameters are chosen so that the constraints on $n_{s}$ and $r$ does not contradict with the current CMB bounds. Whereas $V_{0}$ can be fixed by matching $A_{s}=2.09\times 10^{-9}$ at the pivot. As we have previously mentioned, $n$ and $q$ vary from $1$ to $4$, which results in the $16$ different class of potential. We have analyzed all the possible combinations. In fig. 6.2 we present the results on the $r-n_{s}$ where Slow-Roll I approximation has been utilized. It is evident from fig. 6.2 that almost all the obtained values of $r$ and $n_{s}$ for a combination of potential parameters $n$ and $q$ are well inside the constraints imposed by the Planck 18. Explicit numerical values of the inflationary observables can be found in table 6.2. A comparison between the two slow-roll approaches for a specific subclass of the PLP model is being given in the fig. 6.3.
@@ VI.2 Inflationary Observables in Slow-Roll II
$N$ $r$ $n_{s}$ $r$ $n_{s}$ $r$ $n_{s}$ $r$ $n_{s}$
$50$ 0.0178204    0.971075 0.00680496   0.964019 0.0050969   0.961195 0.00487111   0.960743
$60$ 0.0139747   0.976535 0.00492388   0.970269 0.003579   0.967613 0.00340481   0.967189
@@
Figure 6.3: The comparison between the Slow Roll I and Slow Roll II for the PLP potential with $q=1$ and $n=1-4$. The legends are self-explanatory.

### v3

@@ V.2 New approximation II
The second way to get $Q$ following [1]
@@
Slow roll parameters in approximation-II can be written as:
@@
Using 5.30, one can obtain the remaining slow roll parameters($\varepsilon_{2},\delta_{2}$). The relation between the number of e-folds and the field can be written as [1]:
@@ VI Inflationary analysis
In the subsequent sections, we will provide the calculation of inflationary observables for PLP in the new slow roll approximations. For our analysis, we choose the following form of the EGB coupling [90, 91, 92]
@@
where $\xi_{1}$, $\xi_{2}$ are the coupling constants and $V_{0}$ is the scale of inflation, which we fix from the scalar power spectrum. Substituting Eq. (3.18) and Eq. (6.37) in Eq. (5.26), we can derive the explicit form of the effective potential:
@@
In our analysis, we take $U_{0}=\frac{1}{2}$ and $M_{P}=1$.
VI.1 Inflationary Observables for $\xi(\phi)=0$
In this section, we will present the analysis of the inflationary observables in the absence of an EGB coupling. Inserting $\xi(\phi)=0$ in Eqs. (2.3, 2.4, and 2.5) we recover equations of standard cold inflation without any contribution from the modified gravity. Following the prescription mentioned in the section (IV) we can obtain the slow roll parameters and inflationary observables. In Fig. 6.1 and Table 6.1 we have presented the analysis of $n_{s}$ and $r$ for different values of $n$ and $q$ for $\xi(\phi)=0$ case.
@@ VI.1 Inflationary Observables for $\xi(\phi)=0$
$\Delta N$ $r$ $n_{s}$ $r$ $n_{s}$ $r$ $n_{s}$ $r$ $n_{s}$
5050 0.0251299    0.970275 0.00819738   0.968738 0.0032   0.967405 0.00151532   0.966337
6060 0.0197209   0.975374 0.00622304   0.974041 0.00238632   0.972889 0.00111689   0.971979
Table 6.1: Table for $r-n_{s}$ for different values of potential parameters when $\xi(\phi)=0$
Figure 6.1: Plots of $r-n_{s}$ obtained by using the standard Slow-Roll approximation with $\xi(\phi)=0$ for different values of $n$ and $q$. The red line denotes $n=1$, the dashed green line is for $n=2$, the dotted blue line is for $n=3$, and the black dot-dashed line signifies $n=4$. The small and large dots of different colors signify $\Delta N=50$ and $\Delta N=60$ respectively. $\Delta N$ denotes the number of e-folds from the end of inflation to the onset of inflation ($\Delta N=N_{end}-N_{onset}$). The light blue shaded area represents the $2-\sigma$ bounds, while the light green shaded area represents the $1-\sigma$ bounds on $n_{s}$ from Planck’18 [14].
VI.2 Numerical Calculation of Inflationary Observables
In this section, we present the numerical calculation of the inflationary observables. We obtain the numerical solutions for $\phi,\chi$ and $Q$ in terms of the number of e-folds ($N$) by solving the system  (2.6). To obtain the solution, we need to solve the system  (2.6) numerically, which requires the initial condition to start with. We choose a value of the initial $\phi$>0 and $\chi<0$, for $\chi<0$, the inflationary trajectory is an attractor. Inserting the value of $\phi$ and $\chi$ into the Eq. 2.7 we compute the initial value of Q, after obtaining the required initial conditions we can easily solve the system(2.6) by evolving the system for a range of e-folds ($N$). Once we obtained the numerical solutions, it is straightforward to compute the slow roll parameters using Eq. (4.19) and (4.20). Inserting the slow roll parameters in Eq. (4.23) and (4.24), we calculate the scalar spectral index ($n_{s}$) and tensor to scalar ratio ($r$). The values of $n_{s}$ and $r$ obtained here are exact without using any approximations. (see Fig. 6.2). The exact numerical values are in Table (6.2). We take $\xi_{1}=10$ and $\xi_{2}=0.4$, these values of the coupling parameters are chosen so that the constraints on $n_{s}$ and $r$ do not contradict with the current CMB bounds.
Figure 6.2: Plots of $r-n_{s}$ obtained by solving system ( 2.6) numerically for different values of $n$ and $q$. The red dot denotes the $n=1$, the green dot is for $n=2$, the blue dot is for $n=3$, and the black dot signifies $n=4$. The small and large dots of different colors signify the $\Delta N=50$ and $\Delta N=60$, respectively. $\Delta N$ denotes the number of e-folds from the end of inflation to the onset of inflation ($\Delta N=N_{end}-N_{onset}$). The light blue shaded area is for the $2-\sigma$ bounds, while the light green shaded area represents the $1-\sigma$ bounds on $n_{s}$ from Planck’18 [14]
Figure 6.3: Plots of $r-n_{s}$ obtained by using Slow-Roll Approximation I for different values of $n$ and $q$. The red line denotes $n=1$, the dashed green line is for $n=2$, the dotted blue line is for $n=3$, and the black dot-dashed line signifies $n=4$. The small and large dots of different colors signify $\Delta N=50$ and $\Delta N=60$ respectively. $\Delta N$ denotes the number of e-folds from the end of inflation to the onset of inflation ($\Delta N=N_{end}-N_{onset}$). The light blue shaded area represents the $2-\sigma$ bounds, while the light green shaded area represents the $1-\sigma$ bounds on $n_{s}$ from Planck’18 [14]
@@ VI.2 Numerical Calculation of Inflationary Observables
$\Delta N$ $r$ $n_{s}$ $r$ $n_{s}$ $r$ $n_{s}$ $r$ $n_{s}$
5050 0.0198034    0.968575 0.00798492   0.96082 0.00611248   0.957782 0.00588756   0.957205
6060 0.0152318   0.974837 0.00562672   0.968004 0.00417223   0.965191 0.00399504   0.964688
Table 6.2: Table for $r-n_{s}$ for different values of potential parameters by solving system ( 2.6) numerically for different values of $n$ and $q$
VI.3 Inflationary Observables in Slow-Roll I
Figure 6.4: Plots of $r-n_{s}$ obtained by using Slow-Roll Approximation II for different values of $n$ and $q$. The red line denotes $n=1$, the dashed green line is for $n=2$, the dotted blue line is for $n=3$, and the black dot-dashed line signifies $n=4$. The small and large dots of different colors signify $\Delta N=50$ and $\Delta N=60$, respectively. $\Delta N$ denotes the number of e-folds from the end of inflation to the onset of inflation ($\Delta N=N_{end}-N_{onset}$). The light blue shaded area represents the $2-\sigma$ bounds, while the light green shaded area represents the $1-\sigma$ bounds on $n_{s}$ from Planck’18 [14]
Of course, the results look cumbersome, to say the least, but one should appreciate the fact that using these techniques, one can still obtain analytical solutions.
Now, we will discuss the PLP inflation in slow-roll approximation I. To calculate the inflationary observables, first we calculate the slow roll parameters, using Eq.( 5.29) and Eq.( 5.28)we compute $\varepsilon_{1},\delta_{1}$. It is then straightforward to obtain $\varepsilon_{2}$ and $\delta_{2}$ using Eq.( 5.30). When studying inflation, another important quantity is the number of e-folds $N$. Thus, from Eq.( 5.31), one can write $N$ for a given $\phi$. To solve Eq.( 5.31), we followed the same prescription as mentioned in [1], and we have taken $N=60$ at the end of inflation, whereas $N=0$ corresponds to the onset of inflation.
We set the initial field value for solving Eq.( 5.31) at the end of inflation, which can be calculated from $\varepsilon_{1}=1$. For a particular value of n=(14)n=(1-4) and q=(14)q=(1-4) with $\xi_{1}=10,\xi_{2}=0.4$ we can calculate the field value where $\varepsilon=1$ and we use this as the end of inflation ($\phi_{e}$). We repeat this for all the possible combinations of $n$ and $q$ after acquiring the values of the $\phi_{e}$ for combinations of $n$ and $q$. We use $\phi_{e}$ as the initial condition to solve Eq.(5.31). It is difficult to solve Eq.(5.31) analytically, so we then employ numerical techniques. Equipped with all the preliminary equations in hand, one can finally compute the scalar spectral index ($n_{s}$), the tensor to the scalar ratio ($r$), and the amplitude of scalar power spectrum ($A_{s}$) using Eq.(4.23),(4.24) and Eq.(4.25) respectively. We keep the same values of coupling parameters $\xi_{1}=10$ and $\xi_{2}=0.4$. Whereas, $V_{0}$ can be fixed by matching $A_{s}=2.09\times 10^{-9}$ at the pivot scale.
As we have previously mentioned $n$ and $q$ vary from $1$ to $4$, which results in the $16$ different classes of potential. We have analyzed all the possible combinations. In fig. 6.3 we present the results on the $r-n_{s}$ where Slow-Roll I approximation has been utilized. It is evident from fig. 6.3 that almost all the obtained values of $r$ and $n_{s}$ for combination of potential parameters $n$ and $q$ are well inside the constraints imposed by the Planck 18. Explicit numerical values of the inflationary observables can be found in Table 6.3.
  $q=1$
  $n=1$   $n=2$   $n=3$   $n=4$
$\Delta N$ $r$ $n_{s}$ $r$ $n_{s}$ $r$ $n_{s}$ $r$ $n_{s}$
5050 0.008604    0.961780 0.006397   0.957867 0.005970   0.9568 0.0059473   0.956564
6060 0.006130   0.969175 0.0043752   0.96545 0.0059704   0.96450 0.00401488   0.96429
  $q=2$
  $n=1$   $n=2$   $n=3$   $n=4$
$\Delta N$ $r$ $n_{s}$ $r$ $n_{s}$ $r$ $n_{s}$ $r$ $n_{s}$
5050 0.0114495    0.96616 0.006667   0.959685 0.0058103   0.957663 0.00580653   0.95712
6060 0.008526   0.972857 0.004638   0.966922 0.00398821   0.965102 0.00393797   0.964671
  $q=3$
  $n=1$   $n=2$   $n=3$   $n=4$
$\Delta N$ $r$ $n_{s}$ $r$ $n_{s}$ $r$ $n_{s}$ $r$ $n_{s}$
5050 0.0149482    0.968519 0.007059   0.960959 0.00581401   0.958181 0.00572878   0.957447
6060 0.0114241   0.974752 0.004972   0.967989 0.00398117   0.965487 0.00389589   0.964897
  $q=4$
  $n=1$   $n=2$   $n=3$   $n=4$
$\Delta N$ $r$ $n_{s}$ $r$ $n_{s}$ $r$ $n_{s}$ $r$ $n_{s}$
5050 0.0188761    0.969662 0.00750292   0.961960 0.0058097   0.958572 0.0056763   0.956782
6060 0.0146428   0.975524 0.00533771   0.968838 0.003992   0.965786 0.00386787   0.965061
VI.4 Inflationary Observables in Slow-Roll II
Using the obtained expressions of $\delta_{1}(\phi)$ and $\varepsilon_{1}(\phi)$ and Eq.( 5.30), one can now calculate parameters $\varepsilon_{2}(\phi)$ and $\delta_{2}(\phi)$. One can again get the analytic expressions for the slow-roll parameters using the new approximation-II as given in A.2. After this, one can, in the same spirit, obtain $n_{s}(\phi)$, $r(\phi)$, and $A_{s}(\phi)$ using Eq. (4.23)–(4.25).
In this section, we will discuss the PLP inflation in slow-roll approximation II. We employ the same method to calculate the inflationary observables as we did in Slow Roll Approximation I. First, we calculate the slow roll parameters ($\varepsilon_{1},\varepsilon_{2},\delta_{1},\delta_{2}$), using Eq.( 5.34) and (5.33), and we compute $\varepsilon_{1},\delta_{1}$. After this, it is straightforward to obtain $\varepsilon_{2}$ and $\delta_{2}$ using Eq.( 5.30). Using Eq.( 5.36), we calculate the number of e-folds $N$ by adopting the same technique as we did in the previous subsection (Slow-Roll Approximation I). Equipped with all the preliminaries, we can finally compute the scalar spectral index ($n_{s}$), tensor to scalar ratio ($r$), and amplitude of scalar power spectrum ($A_{s}$) using Eq.( 4.23), (4.24), and Eq.( 4.25), respectively. Here, we also take $\xi_{1}=10$ and $\xi_{2}=0.4$; these values of the coupling parameters are chosen so that the constraints on $n_{s}$ and $r$ do not contradict the current CMB bounds. Whereas $V_{0}$ can be fixed by matching $A_{s}=2.09\times 10^{-9}$ at the pivot. As we have previously mentioned, $n$ and $q$ vary from $1$ to $4$, which results in the $16$ different classes of potential. We have analyzed all the possible combinations. In fig. 6.4, we present the results on $r-n_{s}$ where the Slow-Roll I approximation has been utilized. It is evident from fig. (6.4) that almost all the obtained values of $r$ and $n_{s}$ for a combination of potential parameters $n$ and $q$ are well within the constraints imposed by the Planck’18. Explicit numerical values of the inflationary observables can be found in Table (6.4). A comparison between the two slow-roll approaches for a specific subclass of the PLP model is given in the fig. (6.5).
  $q=1$
  $n=1$   $n=2$   $n=3$   $n=4$
$\Delta N$ $r$ $n_{s}$ $r$ $n_{s}$ $r$ $n_{s}$ $r$ $n_{s}$
5050 0.00729363    0.96551 0.0052205   0.9611672 0.0049137   0.960799 0.00485738   0.96068
6060 0.00535748   0.971735 0.0037563   0.96808 0.0034367   0.967247 0.00339305   0.967135
  $q=2$
  $n=1$   $n=2$   $n=3$   $n=4$
$\Delta N$ $r$ $n_{s}$ $r$ $n_{s}$ $r$ $n_{s}$ $r$ $n_{s}$
5050 0.0103358    0.968584 0.00580519   0.962561 0.00497474   0.960935 0.00486187   0.960702
6060 0.0078487   0.974514 0.00413578   0.968917 0.00348436   0.967372 0.00339693   0.967153
  $q=3$
  $n=1$   $n=2$   $n=3$   $n=4$
$\Delta N$ $r$ $n_{s}$ $r$ $n_{s}$ $r$ $n_{s}$ $r$ $n_{s}$
5050 0.0138884    0.970306 0.00630019   0.963339 0.00503582   0.961067 0.00486645   0.960723
6060 0.0107645   0.975965 0.0045256   0.969642 0.00353203   0.967494 0.00340085   0.967171
  $q=4$
  $n=1$   $n=2$   $n=3$   $n=4$
$\Delta N$ $r$ $n_{s}$ $r$ $n_{s}$ $r$ $n_{s}$ $r$ $n_{s}$
5050 0.0178204    0.971075 0.00680496   0.964019 0.0050969   0.961195 0.00487111   0.960743
6060 0.0139747   0.976535 0.00492388   0.970269 0.003579   0.967613 0.00340481   0.967189
@@ VI.4 Inflationary Observables in Slow-Roll II
Figure 6.5: The comparison between the Slow Roll I and Slow Roll II for the PLP potential with $q=1$ and n=14n=1-4. The legends are self-explanatory.